# Inference-Time Alignment for Diffusion Models via Doob’s Matching

### v1

@@ Abstract
Inference-time alignment for diffusion models aims to adapt a pre-trained diffusion model toward a target distribution without retraining the base score network, thereby preserving the generative capacity of the base model while enforcing desired properties at the inference time. A central mechanism for achieving such alignment is guidance, which modifies the sampling dynamics through an additional drift term. In this work, we introduce Doob’s matching, a novel framework for guidance estimation grounded in Doob’s $h$-transform. Our approach formulates guidance as the gradient of logarithm of an underlying Doob’s $h$-function and employs gradient-penalized regression to simultaneously estimate both the $h$-function and its gradient, resulting in a consistent estimator of the guidance. Theoretically, we establish non-asymptotic convergence rates for the estimated guidance. Moreover, we analyze the resulting controllable diffusion processes and prove non-asymptotic convergence guarantees for the generated distributions in the 2-Wasserstein distance.
@@ 1  Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2021) have emerged as powerful generative tools for sampling from data distributions, achieving remarkable success across diverse domains, including text-to-image and text-to-video generation (Ramesh et al., 2021), Bayesian inverse problems (Chung et al., 2023; Song et al., 2023; Chen et al., 2025; Chang et al., 2025b), and scientific applications (Bao et al., 2024; Li et al., 2025; Si and Chen, 2025; Ding et al., 2024b; Uehara et al., 2025a). Recent years have witnessed the development of large-scale diffusion models pre-trained on vast datasets. Despite the robust capabilities of these base models in capturing the training distribution, the target distributions of downstream generative tasks rarely align perfectly with this base distribution. For example, in conditional generative learning (Dhariwal and Nichol, 2021; Ho and Salimans, 2021), base diffusion models generate samples from a mixture of distributions, whereas downstream tasks require sampling from specific constituent distributions. Similarly, in posterior sampling contexts (Chung et al., 2023; Song et al., 2023; Chen et al., 2025; Chang et al., 2025b; Purohit et al., 2025; Martin et al., 2025), the base distribution serves as the prior, while the target is the posterior defined by tilting the prior with a measurement likelihood. Furthermore, downstream tasks frequently impose additional constraints, such as human preferences or safety considerations (Domingo-Enrich et al., 2024, 2025; Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025), which must be satisfied without compromising the model’s generative quality.
@@
To bridge the gap between base and target distributions, researchers have proposed numerous alignment methods (Xu et al., 2023; Lee et al., 2023b; Fan et al., 2023; Clark et al., 2024; Domingo-Enrich et al., 2024, 2025; Uehara et al., 2024b, 2025b). These strategies generally fall into two categories: fine-tuning and inference-time alignment. Fine-tuning approaches involve retraining the base score network via supervised learning (Lee et al., 2023b), reinforcement learning (Fan et al., 2023; Black et al., 2024; Clark et al., 2024; Uehara et al., 2024c), or classifier-free fine-tuning (Ho and Salimans, 2021; Zhang et al., 2023; Yuan et al., 2023). Despite its conceptual simplicity, fine-tuning presents significant limitations. First, it often requires a substantial collection of high-quality samples from the target distribution, which may be unavailable in practical scenarios like posterior sampling. Second, the computational cost of retraining score networks can be prohibitive, particularly for large-scale models with billions of parameters (Uehara et al., 2025b). Third, fine-tuning is vulnerable to over-optimization (Gao et al., 2023; Rafailov et al., 2024; Kim et al., 2025), where the network overfits to limited target samples or preferences, causing it to “forget” the valuable prior information encoded in the base model. This degradation undermines the fundamental advantage of leveraging pre-trained models.
@@
In contrast, inference-time alignment (Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025; Pachebat et al., 2025) eliminates the need to retrain the base diffusion model. These methods offer substantial computational advantages and preserve the generative capacity of the underlying model. The core technique is guidance (Jiao et al., 2025), which incorporates target information by introducing an additional drift term to the base diffusion model. Within the framework of Doob’s $h$-transform (Rogers and Williams, 2000; Särkkä and Solin, 2019; Chewi, 2025), the score function for the target tilted distribution decomposes into the sum of the base score and a guidance term, where the guidance is defined as the gradient of the log-Doob’s $h$-function (Heng et al., 2024; Tang and Xu, 2024; Denker et al., 2024, 2025). This relationship has also been investigated through the lens of classifier guidance (Dhariwal and Nichol, 2021), stochastic optimal control (Han et al., 2024; Tang and Zhou, 2025), and Bayes’ rule (Chung et al., 2023; Song et al., 2023). The primary challenge lies in accurately estimating this guidance.
@@
To address these limitations, we introduce Doob’s $h$-matching, a novel framework for provable guidance estimation in inference-time alignment. We propose a gradient-penalized regression method that simultaneously estimates Doob’s $h$-function and its gradient, yielding a consistent estimator of the guidance. When combined with the pre-trained base score, our method enables efficient sampling from the target distribution without the need for computationally expensive fine-tuning or access to target distribution samples.
@@ 1.1  Contributions
    We introduce Doob’s matching, a novel guidance estimation framework for controllable diffusion models grounded in Doob’s $h$-transform. The Doob $h$-function is estimated via a least-squares regression approach augmented with a gradient penalty, and the gradient of the logarithm of the resulting $h$-function estimator yields an estimator for the Doob’s guidance. Additionally, this method is derivative-free, meaning it does not require access to the gradient of the weight function between the target tilted distribution and the base distribution.
@@
    We establish non-asymptotic convergence rates for Doob’s matching, showing that the proposed method guarantees convergence of both the $h$-function estimator and its gradient. These results provide rigorous theoretical guarantees for Doob’s guidance estimation (Theorem 4.3). Moreover, we derive non-asymptotic convergence rates for the induced controllable diffusion models, thereby establishing rigorous guarantees for the generated distributions in the $2$-Wasserstein distance (Theorem 4.5).
@@ 2.1  Forward and time-reversal process
We consider the diffusion model for a base distribution $p_{0}$. The forward process of the base diffusion model is defined by the Ornstein–Uhlenbeck process:
@@
$\mathbf{X}_{0}^{\leftarrow}$ $\sim p_{T},$
where the marginal density of $\mathbf{X}_{t}^{\leftarrow}$ coincides $p_{T-t}$ for each $t\in(0,T)$.
@@ 2.2  Path measure and filtration
2.3  Training phase: base score matching
In generative learning, the exact base score $\nabla\log p_{t}$ in the time-reversal process (2.4) is intractable. One can estimate the base score using samples from the base density $p_{0}$ via standard techniques such as implicit score matching (Hyvärinen, 2005), sliced score matching (Song et al., 2020), and denoising score matching (Vincent, 2011). Let $\widehat{\mathbf{s}}:(0,T)\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ denote an estimator for the prior score, that is,
@@ 2.3  Training phase: base score matching
for a small tolerence $\varepsilon\ll 1$. Considerable research has established theoretical guarantees for this score estimation (Tang and Yang, 2024; Oko et al., 2023; Fu et al., 2024; Ding et al., 2024a), leveraging standard techniques from non-parametric regression with deep neural networks (Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Jiao et al., 2023).
@@ 2.4  Inference phase: sampling
Given a base score estimator $\widehat{\mathbf{s}}$ in (2.5), the inference phase of diffusion models aims to generate samples by simulating the time-reversal process with estimated score. Since the explicit solution of the time-reversal process is intractable, we employ an exponential integrator (Hochbruck and Ostermann, 2005, 2010; Lu et al., 2022a; Zhang and Chen, 2023). This approach is well-suited for solving the time-reversal process due to the semi-linearity of the drift term of the SDE in (2.4).
@@
3  Controllable Diffusion Models via Doob’s Matching
In this section, we propose controllable diffusion models for sampling from a target distribution, defined as the base distribution tilted by a weight function. We utilize the theory of measure change for diffusion processes on the filtered space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$.
@@ 3.1  Problem setup
We assume access to a pre-trained base diffusion model (2.6) that generates samples approximately from the base distribution $p_{0}$. We aim to sample from a tilted distribution $q_{0}$, defined by reweighting the base distribution with a known weight function $w:{\mathbb{R}}^{d}\to{\mathbb{R}}_{\geq 0}$:
@@
Our goal is to derive a new diffusion process that generates samples from the tilted distribution $q_{0}$ directly by introducing a drift correction to the pre-trained base diffusion model (2.6).
@@
Bayesian inverse problems play a critical role in scientific computing (Stuart, 2010; Kantas et al., 2014; Ding et al., 2024b), image science (Chung et al., 2023; Mardani et al., 2024; Purohit et al., 2025; Chang et al., 2025b), and medical imaging (Song et al., 2022). In Bayesian inverse problems, we aim to recover an unknown signal $\mathbf{X}_{0}\in{\mathbb{R}}^{d}$ from noisy measurements $\mathbf{Y}\in{\mathbb{R}}^{m}$, which are linked by the following measurement model:
@@
where $\mathcal{A}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{m}$ is a known measurement operator, and $\mathbf{n}\in{\mathbb{R}}^{m}$ represents a measurement noise with a known distribution. The Bayesian approach incorporates prior knowledge about $\mathbf{X}_{0}$ in the form of a prior distribution $p_{0}$. Given observed measurements $\mathbf{Y}=\mathbf{y}$, the goal of Bayesian inverse problems is to sample from the posterior distribution:
@@
In Bayesian inverse problems, one typically has a base diffusion model pre-trained on the prior distribution, and aims to sample from the posterior distribution (3.3) without retraining the base model.
@@
In the reward-guided generation (Domingo-Enrich et al., 2025; Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025), human preferences and constraints can be encoded into a reward function $r:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$. For instance, in text-to-image generation, the reward function $r$ quantifies how well the generated data aligns with the input prompt. In practice, such reward function can be learned from the human feedback or preference data (Stiennon et al., 2020; Ouyang et al., 2022; Lee et al., 2023b). For the sake of simplicity, we assume throughout this work that the reward function has already been given. A naive approach to reward-guided generation is to maximize the expected reward $\max_{\pi\in{\mathcal{P}}}{\mathbb{E}}_{\pi}[r]$, where $\mathcal{P}$} is the set of probability measures on ${\mathbb{R}}^{d}$. However, solely maximizing the expected reward may lead to over-optimization and degenerate solutions (Kim et al., 2025). To mitigate this, entropy regularization (Uehara et al., 2024a; Tang, 2024) is incorporated into the objective functional, yielding the following optimization problem:
@@
where $\alpha>0$ is a regularization parameter, and $p_{0}$ is the density of a reference distribution, i.e., the distribution corresponding to the pre-trained base diffusion model. This objective comprises two components: the expected reward, which captures human preferences, and the KL-divergence term, which prevents the distribution from deviating excessively from the base diffusion model. The closed-form solution to this optimization problem (3.5) is given by (Rafailov et al., 2023):
@@
where $Z$ is the partition function to ensure $q_{0}$ is a valid probability density. In reward-guided generation, the central objective is to incorporate preferences exclusively during the inference phase, thereby avoiding the substantial computational cost of retraining the large-scale base diffusion model.
@@
Diffusion models have achieved remarkable success in image generation. However, their performance critically depends on the availability of large-scale training data. In the scenarios of few-shot generation, training diffusion models solely on limited samples typically results in poor generative performance. To retain strong generative capability under data scarcity, a common strategy is to transfer expressive diffusion models pre-trained on large datasets to the target domain (Ouyang et al., 2024; Wang et al., 2024; Zhong et al., 2025; Bahram et al., 2026). Formally, transfer learning aims to adapt a model pre-trained on a large-scale source distribution to a target distribution of much smaller size and diversity of samples. However, directly fine-tuning a large pre-trained diffusion model using only limited target samples often leads to severe overfitting (Wang et al., 2024). To address this issue, transfer learning approaches for diffusion models typically train a lightweight guidance network on the limited target data and combine it with the pre-trained base score network, yielding a modified diffusion model capable of sampling from the target distribution (Ouyang et al., 2024; Zhong et al., 2025; Bahram et al., 2026). Concretely, Ouyang et al. (2024) estimates the density ratio between the target and source distributions,
@@
using limited samples from the target dsitribution. For simplicity, we assume throughout this work that the density ratio is known. Under this assumption, the objective of transfer learning for diffusion models is to sample from the target distribution $q_{0}$ by leveraging the pre-trained base diffusion model together with the estimated density ratio.

### v2

# Inference-Time Alignment for Diffusion Models via Variationally Stable Doob’s Matching

@@ Abstract
Inference-time alignment for diffusion models aims to adapt a pre-trained reference diffusion model toward a target distribution without retraining the reference score network, thereby preserving the generative capacity of the reference model while enforcing desired properties at the inference time. A central mechanism for achieving such alignment is guidance, which modifies the sampling dynamics through an additional drift term. In this work, we introduce variationally stable Doob’s matching, a novel framework for provable guidance estimation grounded in Doob’s $h$-transform. Our approach formulates guidance as the gradient of logarithm of an underlying Doob’s $h$-function and employs gradient-regularized regression to simultaneously estimate both the $h$-function and its gradient, resulting in a consistent estimator of the guidance. Theoretically, we establish non-asymptotic convergence rates for the estimated guidance. Moreover, we analyze the resulting controllable diffusion processes and prove non-asymptotic convergence guarantees for the generated distributions in the 2-Wasserstein distance. Finally, we show that variationally stable guidance estimators are adaptive to unknown low dimensionality, effectively mitigating the curse of dimensionality under low-dimensional subspace assumptions.
@@ 1  Introduction
Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song and Ermon, 2019; Song et al., 2021) have emerged as powerful generative tools for sampling from data distributions, achieving remarkable success across diverse domains, including text-to-image and text-to-video generation (Ramesh et al., 2021), Bayesian inverse problems (Chung et al., 2023; Song et al., 2023; Chen et al., 2025; Chang et al., 2025b), and scientific applications (Bao et al., 2024; Li et al., 2025; Si and Chen, 2025; Ding et al., 2024; Uehara et al., 2025a). Recent years have witnessed the development of large-scale diffusion models pre-trained on vast datasets. Despite the robust capabilities of these reference models in capturing the training distribution, the target distributions of downstream generative tasks rarely align perfectly with this reference distribution. For example, in conditional generative learning (Dhariwal and Nichol, 2021; Ho and Salimans, 2021), reference diffusion models generate samples from a mixture of distributions, whereas downstream tasks require sampling from specific constituent distributions. Similarly, in posterior sampling contexts (Chung et al., 2023; Song et al., 2023; Chen et al., 2025; Chang et al., 2025b; Purohit et al., 2025; Martin et al., 2025), the reference distribution serves as the prior, while the target is the posterior defined by tilting the prior with a measurement likelihood. Furthermore, downstream tasks frequently impose additional constraints, such as human preferences or safety considerations (Domingo-Enrich et al., 2024, 2025; Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025), which must be satisfied without compromising the model’s generative quality.
@@
To bridge the gap between reference and target distributions, researchers have proposed numerous alignment methods (Xu et al., 2023; Lee et al., 2023b; Fan et al., 2023; Clark et al., 2024; Domingo-Enrich et al., 2024, 2025; Uehara et al., 2024b, 2025b). These strategies generally fall into two categories: fine-tuning and inference-time alignment. Fine-tuning approaches involve retraining the reference score network via supervised learning (Lee et al., 2023b), reinforcement learning (Fan et al., 2023; Black et al., 2024; Clark et al., 2024; Uehara et al., 2024c), or classifier-free fine-tuning (Ho and Salimans, 2021; Zhang et al., 2023; Yuan et al., 2023). Despite its conceptual simplicity, fine-tuning presents significant limitations. First, it often requires a substantial collection of high-quality samples from the target distribution, which may be unavailable in practical scenarios like posterior sampling. Second, the computational cost of retraining score networks can be prohibitive, particularly for large-scale models with billions of parameters (Uehara et al., 2025b). Third, fine-tuning is vulnerable to over-optimization (Gao et al., 2023; Rafailov et al., 2024; Kim et al., 2025), where the network overfits to limited target samples or preferences, causing it to “forget” the valuable prior information encoded in the reference model. This degradation undermines the fundamental advantage of leveraging pre-trained models.
@@
In contrast, inference-time alignment (Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025; Pachebat et al., 2025) eliminates the need to retrain the reference diffusion model. These methods offer substantial computational advantages and preserve the generative capacity of the underlying model. The core technique is guidance (Jiao et al., 2025), which incorporates target information by introducing an additional drift term to the reference diffusion model. Within the framework of Doob’s $h$-transform (Rogers and Williams, 2000; Särkkä and Solin, 2019; Chewi, 2025), the score function for the target tilted distribution decomposes into the sum of the reference score and a guidance term, where the guidance is defined as the gradient of the log-Doob’s $h$-function (Heng et al., 2024; Tang and Xu, 2024; Denker et al., 2024, 2025). This relationship has also been investigated through the lens of classifier guidance (Dhariwal and Nichol, 2021), stochastic optimal control (Han et al., 2024; Tang and Zhou, 2025), and Bayes’ rule (Chung et al., 2023; Song et al., 2023). The primary challenge lies in accurately estimating this guidance.
@@
To address these limitations, we introduce variationally stable Doob’s $h$-matching, a novel framework for provable guidance estimation in inference-time alignment. We propose a gradient-regularized regression method that simultaneously estimates Doob’s $h$-function and its gradient, yielding a consistent estimator of the guidance. When combined with the pre-trained reference score, our method enables efficient sampling from the target distribution without the need for computationally expensive fine-tuning or access to target distribution samples.
@@ 1.1  Contributions
    We introduce variationally stable Doob’s matching, a novel guidance estimation framework for controllable diffusion models grounded in Doob’s $h$-transform. The Doob $h$-function is estimated via a least-squares regression approach augmented with a gradient regularization, and the plug-in gradient of the logarithm of the resulting $h$-function estimator yields an estimator for the Doob’s guidance. Additionally, this method is derivative-free, meaning it does not require access to the gradient of the weight function between the target tilted distribution and the reference distribution.
@@
    We establish non-asymptotic convergence rates for variationally stable Doob’s matching, showing that the proposed method guarantees convergence of both the $h$-function estimator and its gradient. These results provide rigorous theoretical guarantees for Doob’s guidance estimation (Theorem 5.3). Moreover, we derive non-asymptotic convergence rates for the induced controllable diffusion models, thereby establishing rigorous guarantees for the generated distributions in the $2$-Wasserstein distance (Theorem 5.6). Additionally, we obtain convergence rates that depend only on the intrinsic dimension, thereby mitigating the curse of dimensionality under low-dimensional subspace assumptions (Theorem 5.9 and Corollary 5.10).
@@ 2.1  Forward and time-reversal process
We consider the diffusion model for a reference distribution $p_{0}$. The forward process of the reference diffusion model is defined by the Ornstein–Uhlenbeck process:
@@
$\mathbf{X}_{0}^{\leftarrow}$ $\sim p_{T}.$
It has been established that the path measure of the time-reversal process $(\mathbf{X}_{t}^{\leftarrow})_{0\leq t\leq T}$ corresponds exactly to the reverse of the forward process $(\mathbf{X}_{t})_{0\leq t\leq T}$ (Anderson, 1982).
@@ 2.2  Path measure and filtration
2.3  Training phase: score matching
In generative learning, the exact reference score $\nabla\log p_{t}$ in the time-reversal process (2.4) is intractable. One can estimate the reference score using samples from the reference density $p_{0}$ via standard techniques such as implicit score matching (Hyvärinen, 2005), sliced score matching (Song et al., 2020), and denoising score matching (Vincent, 2011). Let $\widehat{\mathbf{s}}:(0,T)\times{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ denote an estimator for the prior score, that is,
@@ 2.3  Training phase: score matching
for a small tolerence $\varepsilon\ll 1$. Considerable research has established theoretical guarantees for this score estimation (Tang and Yang, 2024; Oko et al., 2023; Fu et al., 2024; Ding et al., 2025a), leveraging standard techniques from non-parametric regression with deep neural networks (Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021; Jiao et al., 2023).
@@ 2.4  Inference phase: sampling
Given a reference score estimator $\widehat{\mathbf{s}}$ in (2.5), the inference phase of diffusion models aims to generate samples by simulating the time-reversal process with estimated score. Since the explicit solution of the time-reversal process is intractable, we employ an exponential integrator (Hochbruck and Ostermann, 2005, 2010; Lu et al., 2022a; Zhang and Chen, 2023). This approach is well-suited for solving the time-reversal process due to the semi-linearity of the drift term of the SDE in (2.4).
@@
3  Controllable Diffusion Models and Doob’s Transform
In this section, we propose controllable diffusion models for sampling from a target distribution, defined as the reference distribution tilted by a weight function. We utilize the theory of measure change for diffusion processes on the filtered space $(\Omega,{\mathcal{F}},{\mathbb{F}},{\mathbb{P}})$.
@@ 3.1  Problem setup
We assume access to a pre-trained reference diffusion model (2.6) that generates samples approximately from the reference distribution $p_{0}$. We aim to sample from a tilted distribution $q_{0}$, defined by reweighting the reference distribution with a known weight function $w:{\mathbb{R}}^{d}\to{\mathbb{R}}_{\geq 0}$:
@@
Our goal is to derive a new diffusion process that generates samples from the tilted distribution $q_{0}$ directly by introducing a drift correction to the pre-trained reference diffusion model (2.6).
@@
Bayesian inverse problems play a critical role in scientific computing (Stuart, 2010; Kantas et al., 2014; Ding et al., 2024), image science (Chung et al., 2023; Mardani et al., 2024; Purohit et al., 2025; Chang et al., 2025b), and medical imaging (Song et al., 2022). In Bayesian inverse problems, we aim to recover an unknown signal $\mathbf{X}_{0}\in{\mathbb{R}}^{d}$ from noisy measurements $\mathbf{Y}\in{\mathbb{R}}^{m}$, which are linked by the following measurement model:
@@
where $\mathcal{A}:{\mathbb{R}}^{d}\to{\mathbb{R}}^{m}$ is a known measurement operator, and $\mathbf{n}\in{\mathbb{R}}^{m}$ represents a measurement noise with a known distribution. The Bayesian approach incorporates prior knowledge about $\mathbf{X}_{0}$ in the form of a prior distribution $p_{0}$. Given observed measurements $\mathbf{Y}=\mathbf{y}$, the goal of Bayesian inverse problems is to sample from the posterior distribution:
@@
In Bayesian inverse problems, one typically has a reference diffusion model pre-trained on the prior distribution, and aims to sample from the posterior distribution (3.3) without retraining the reference model.
@@
In the reward-guided generation (Domingo-Enrich et al., 2025; Uehara et al., 2025b; Kim et al., 2025; Sabour et al., 2025; Denker et al., 2025; Ren et al., 2025), human preferences and constraints can be encoded into a reward function $r:{\mathbb{R}}^{d}\to{\mathbb{R}}$. For instance, in text-to-image generation, the reward function $r$ quantifies how well the generated data aligns with the input prompt. In practice, such reward function can be learned from the human feedback or preference data (Stiennon et al., 2020; Ouyang et al., 2022; Lee et al., 2023b). For the sake of simplicity, we assume throughout this work that the reward function has already been given. A naive approach to reward-guided generation is to maximize the expected reward $\max_{\pi\in{\mathcal{P}}}{\mathbb{E}}_{\pi}[r]$, where $\mathcal{P}$} is the set of probability measures on ${\mathbb{R}}^{d}$. However, solely maximizing the expected reward may lead to over-optimization and degenerate solutions (Kim et al., 2025). To mitigate this, entropy regularization (Uehara et al., 2024a; Tang and Zhou, 2025) is incorporated into the objective functional, yielding the following optimization problem:
@@
where $\alpha>0$ is a regularization parameter, and $p_{0}$ is the density of a reference distribution, i.e., the distribution corresponding to the pre-trained reference diffusion model. This objective comprises two components: the expected reward, which captures human preferences, and the KL-divergence term, which prevents the distribution from deviating excessively from the reference diffusion model. The closed-form solution to this optimization problem (3.5) is given by (Rafailov et al., 2023):
@@
where $Z$ is the partition function to ensure $q_{0}$ is a valid probability density. In reward-guided generation, the central objective is to incorporate preferences exclusively during the inference phase, thereby avoiding the substantial computational cost of retraining the large-scale reference diffusion model.
@@
Diffusion models have achieved remarkable success in image generation. However, their performance critically depends on the availability of large-scale training data. In the scenarios of few-shot generation, training diffusion models solely on limited samples typically results in poor generative performance. To retain strong generative capability under data scarcity, a common strategy is to transfer expressive diffusion models pre-trained on large datasets to the target domain (Ouyang et al., 2024; Wang et al., 2024; Zhong et al., 2025; Bahram et al., 2026). Formally, transfer learning aims to adapt a model pre-trained on a large-scale source distribution to a target distribution of much smaller size and diversity of samples. However, directly fine-tuning a large pre-trained diffusion model using only limited target samples often leads to severe overfitting (Wang et al., 2024). To address this issue, transfer learning approaches for diffusion models typically train a lightweight guidance network on the limited target data and combine it with the pre-trained reference score network, yielding a modified diffusion model capable of sampling from the target distribution (Ouyang et al., 2024; Zhong et al., 2025; Bahram et al., 2026). Concretely, Ouyang et al. (2024) estimates the density ratio between the target and source distributions,
@@
using limited samples from the target dsitribution. For simplicity, we assume throughout this work that the density ratio is known. Under this assumption, the objective of transfer learning for diffusion models is to sample from the target distribution $q_{0}$ by leveraging the pre-trained reference diffusion model together with the estimated density ratio.